# Anomalous diffusion from hydrodynamic recoupling in particle-hole symmetric fluids

### v1

@@ Abstract
Hydrodynamics dictates how systems relax to equilibrium under generic classical or quantum dynamics landau1987fluid; Chaikin_Lubensky_1995; forster2018hydrodynamic; spohn2012large; liu2018lectures. The predictions of hydrodynamics depend only on the symmetries of the system; long-lived deviations from equilibrium are due to fluctuations of the associated conserved charge densities. Within hydrodynamics, charge fluctuations generically relax diffusively (if momentum is not conserved) or spread ballistically (if it is). In the momentum-conserving case, charge density fluctuations are carried by ballistically propagating sound waves. It has long been appreciated that discrete symmetries can change this picture: for example, charge conjugation (or particle-hole) symmetry can prevent sound waves from coupling to a particular conserved charge PhysRevB.57.8307; Fong2012; Lucas2018; anomalousFCS_spinchain_review. In this situation—which occurs most famously in graphene at charge neutrality crossno2016observation; majumdar2025universality—ballistic energy transport coexists with diffusive charge transport. This symmetry-protected charge diffusion qualitatively differs from the “regular” hydrodynamic diffusion that occurs in the presence of momentum relaxation: as we recently showed pnas.2403327121; PhysRevE.111.015410; Gopalakrishnan2022a, symmetry-protected diffusion (in quasi-1D geometries) leads to anomalous noise signatures, such as non-gaussian full counting statistics Levitov1996ElectronCS.
@@
We will describe these terms in detail below. The contribution $D_{\rm Fick}$ is regular as a function of $\gamma_{e},\gamma_{p}$, with (generically) a nonzero value when $\gamma_{e}=\gamma_{p}=0$. However, at the symmetric point, there is an additional convective contribution $D_{\rm conv}$ from the ballistic sound waves to the charge diffusion constant Gopalakrishnan2018; Medenjak2020; pnas.2403327121; PhysRevE.111.015410; PhysRevLett.134.187101. In the momentum and energy-breaking case, the diffusion constant instead scales as
@@
with $\ldots$ representing corrections that are analytic in $\gamma_{e},\gamma_{p}$ (and disappearing as $\gamma_{e},\gamma_{p}\to 0$suppmat. This result has important implications for recent computational approaches (on both classical and quantum computers) that aim to compute diffusion constants by zero-noise extrapolation temme2016error; li2017efficient; RevModPhys.95.045005; PhysRevB.105.245101; xct1-7kf2: the system we are considering presents an explicit example where the diffusion constant converges to some value at low noise, but this converged value is unrelated to the true zero-noise diffusion constant $\mathfrak{D}$. Previous work had established a discontinuity of this type in the fine-tuned case of integrable spin chains 10.1073/pnas.2202823119 subject to dephasing noise; we show here that the conclusion is in fact completely general, relying neither on integrability nor on the noise model.
Another key result of our work is that the diffusion constant diverges when $\gamma_{p}=0$ and $\gamma_{e}\neq 0$: i.e., when only momentum and charge are conserved, and particle-hole symmetry forbids them from coupling linearly to one another. Thus, weak momentum-conserving noise (e.g., global fluctuations of the Hamiltonian parameters) can cause a singular enhancement of charge diffusion, leading to superdiffusive charge transport. The coupled momentum and charge dynamics can be described using a recent hydrodynamic theory m_phi_theory; roy2024universality of coupled Burgers equations. This theory was introduced to describe certain fine-tuned limits of coupled asymmetric exclusion processes; we show how it emerges without fine-tuning.
Three-mode hydrodynamics.—We briefly review the three-mode hydrodynamics of particle-hole symmetric fluids in quasi-one-dimensional geometries pnas.2403327121; PhysRevE.111.015410; PhysRevE.111.024141; yoshimura2025hydrodynamicfluctuationsstochasticcharged. In this work we will be concerned with linear response about equilibrium states at nonzero temperature. We consider systems with three conserved densities: energy $e$, momentum $p$, and charge $n$. These transform under discrete symmetries as follows: $e,n$ are even under time-reversal and spatial reflection, while $p$ is odd under these symmetries; $e,p$ are even under charge conjugation, while $n$ is odd under this symmetry. (The chemical potential conjugate to $n$ is pinned by charge conjugation symmetry.) Given these constraints, we arrive at the following hydrodynamic equations pnas.2403327121:
@@
In these equations, $\eta$ denotes Gaussian white noise whose strength is set by the fluctuation-dissipation theorem, $c$ is the speed of sound, and the coefficients $C_{2},D_{\rm Fick}$ are $O(1)$ nonuniversal constants. For convenience we set $C_{2}=1$. Also, $\ldots$ denotes terms of higher order in the gradient expansion, which can be neglected for the purposes of understanding the hydrodynamics of charge pnas.2403327121. Among these neglected terms is the renormalization of the speed of sound $c$ by equilibrium charge fluctuations. This renormalization can be absorbed into the definition of $c$, so that the equations for $e,p$ form a closed system describing sound waves that that carry energy and momentum (but not charge).
@@
We will now compute this diffusion constant in two steps. First, we will artificially set $D_{\rm Fick}=0$. In this limit, the density obeys the first-order differential equation $\partial_{t}n+\partial_{x}(pn)=0$. This equation can be solved by dividing the density into small packets, each with an evolving position: $n(x,t)=\int dx_{0}\,n(x_{0},0)\,\delta(x-X(x_{0},t))$, with characteristics solving $\partial_{t}X=p(X(x_{0},t),t)$. A crucial simplification is that the ballistic sound waves move parametrically faster than the packets of charge; therefore, one can replace $p(X,t)$ with $p(x_{0},t)$ up to subleading corrections pnas.2403327121. Making this approximation, one arrives at a rigid shift of the charge pattern,
@@
To complete the calculation, we restore Fickian processes, as follows. Consider conditioning on a particular history of sound waves and a particular initial charge configuration. On this history, the charge pattern is rigidly translated by $X(t)$. Beyond this rigid translation, the charge dynamics is due to Fickian processes. If one treats the elements of charge as particles, one can describe these two effects as follows: the particle configuration is randomly rearranged over the scale $\sqrt{D_{\rm Fick}t}$ in addition to being shifted by $\sqrt{D_{\rm conv}t}$. The full diffusion constant comes from convolving these effects pnas.2403327121; yoshimura2025hydrodynamicfluctuationsstochasticcharged, $n(x,t)=\int dx^{\prime}G_{\rm Fick}(x-x^{\prime},t)n_{0}(x^{\prime}-X(t))$: for each fixed $X$, the momentum-space charge structure factor at momentum $k$ is $S_{X}(k,t)=\exp[-D_{\mathrm{Fick}}k^{2}t+ikX(t)]$; averaging this structure factor over the Gaussian variable $X(t)$ gives the final result $S(k,t)=\exp[-(D_{\mathrm{Fick}}+D_{\mathrm{conv}})k^{2}t]$, which yields Eq. (1). Since the convective and Fickian effects can be treated separately, we will focus below on the fate of the convective contribution when relaxation processes are allowed; the Fickian contribution can be restored at the end of the calculation.
@@
where $S_{pp}(k,\omega)$ is the Fourier transform of $\langle p(x,t)p(0,0)\rangle$ (see suppmat). The diffusive broadening of the sound peaks is subleading to the Euler-scale advection in the weak symmetry breaking limit ($\gamma_{e},\gamma_{p}\to 0$). For simplicity, we now set $D_{p}=0$, (see the supplemental materials suppmat for the general case). Fourier transforming Eqs. (6) and solving for $p(k,\omega)$ yields
@@
where $\chi_{e}$ and $\chi_{p}$ are energy and momentum static susceptibilities. This expression has two immediate and striking consequences. First, the diffusion constant along a general ray in $(\gamma_{e},\gamma_{p})$ space converges to a finite ray-dependent value that does not match continuously to $\mathfrak{D}$; thus the zero-noise limit is singular. Second, it suggests that when energy (momentum) remains as a residual conserved quantity, the diffusion constant vanishes (diverges). To gain an intuition for Eq. (9) we now consider these two limiting cases, for which the analysis simplifies.
@@
Energy-conserving limit.—First, we consider setting $\gamma_{e}=0$, for example by considering static disorder instead of noise. In this case, energy evolves diffusively, with a diffusion constant set by $D_{e}\sim 1/\gamma_{p}$. By Lorentz covariance, momentum density is equal to the energy current: consequently, $X(t)$ is the integrated current of a diffusive quantity across a fixed point, and as such scales as $X(t)\sim t^{1/4}$. Thus, the convective contribution to charge transport becomes subdiffusive with dynamical exponent $z=4$ for any nonzero $\gamma_{p}$ when $\gamma_{e}=0$. In this regime, the convective contribution to charge transport is subleading to the Fickian term, and the diffusion constant is simply $D_{\rm Fick}$ as $\gamma_{p}\to 0$. Intuitively, the subdiffusive behavior can be understood as follows (see also PhysRevLett.127.230602; 10.1073/pnas.2202823119). In time $t$, a given packet of charge is repeatedly encountered by the same sound waves as they scatter back and forth. Since the backscattering of sound waves conserved energy, its momentum changes sign at each backscatter, so that successive encounters with a wave give opposite kicks. As a result, each sound wave contributes only an $O(1)$ random kick to the net displacement of a charge packet over time $t$. Since the sound waves themselves diffuse, only $O(t^{1/2})$ independent sound waves cross the origin in time $t$. As a sum of $t^{1/2}$ random kicks, the net displacement grows as $X(t)\sim t^{1/4}$. This limit clearly illustrates the crucial difference between the $\gamma_{p}=0$ and $\gamma_{p}\neq 0$ cases: in the former case each charge packet encounters each sound wave exactly once (by hydrodynamic decoupling), but for any nonzero $\gamma_{p}$ the charge packet encounters each sound wave infinitely many times (in the quasi-1D geometry we are considering).
@@
Eq. (Anomalous diffusion from hydrodynamic recoupling in particle-hole symmetric fluids) describes a “degenerate” limit of two coupled degenerate Burgers equations, recently introduced in Ref. m_phi_theory (see also PhysRevLett.69.929) and explored further in Refs. Roy_2024; roy2025fixedpointsuniversalityclasses (where the equations take “cyclic” form yielding a Gaussian stationary state). Previously considered realizations of Eq. (Anomalous diffusion from hydrodynamic recoupling in particle-hole symmetric fluids) required fine-tuning to achieve the degenerate and cyclic conditions; in the present case, these conditions follow from the symmetries of the problem. Momentum and charge are strongly nonlinearly coupled, and both quantities scale with the Kardar-Parisi-Zhang (KPZ) PhysRevLett.56.889 dynamical exponent $x\sim t^{2/3}$.
@@
In the absence of relaxation processes, $\langle p(x,t)p(x,0)\rangle$ decorrelates rapidly, since sound waves move ballistically. Energy relaxation without momentum relaxation causes positive correlations, $\langle p(x,t)p(0,0)\rangle\sim(\gamma_{e}t)^{-1/2}$ on timescales starting with the first back-scattering event $\sim\gamma_{e}^{-1}$ (since momentum is now a diffusive mode with a large return probability). Conversely, momentum relaxation without energy relaxation causes sound waves to back-scatter, giving rise to anticorrelations in $\langle p(x,t)p(x,0)\rangle\sim-\gamma_{p}^{-1/2}t^{-3/2}$ (since a returning sound wave is carrying the same energy but opposite momentum) PhysRevB.73.035113. The slower of the two relaxation processes acts as a long-time cutoff: beyond that timescale, sound waves cease to be defined suppmat, and $n$ is the only surviving hydrodynamic mode. For example, when energy relaxation is faster, the power-law dependence of $\langle p(x,t)p(0,0)\rangle$ is cut off on a timescale $\sim\gamma_{p}^{-1}$, so the net contribution to the diffusion constant is $\sqrt{\gamma_{p}/\gamma_{e}}$. Even though the current autocorrelation function is always instantaneously close to its value in the absence of relaxation processes, the emergent long timescale gives rise to a singular correction to diffusion (Fig. 1). We dub this phenomenon “hydrodynamic recoupling”, as the sound waves that decoupled in the $\gamma_{e}=\gamma_{n}=0$ limit “recouple” in a singular way acting as a non-Markovian bath when $\gamma_{e},\gamma_{n}>0$.
@@
On the diagonal ray $\gamma_{e}=\gamma_{p}$ backscattering of sound waves is equally likely to reverse the wave’s momentum as not, so that consecutive collisions with a charge element provide uncorrelated kicks. Therefore, charge diffusion in the weak dissipation limit along this special ray corresponds to the fully conserved case—in both cases the kicks contributing to $X(t)$ are effectively Markovian suppmat—fixing the coefficient of $\sqrt{\gamma_{e}/\gamma_{p}}$ to $D_{\rm conv}$ in Eq. (2). Finally we remark that the scaling form (9) will be modified for $\gamma_{p}/\gamma_{e}\ll 1$, because the nonlinearities in Eq. (Anomalous diffusion from hydrodynamic recoupling in particle-hole symmetric fluids) must be included in the solution of Eq. (6).
@@
Charge transfer statistics.—While our discussion of charge transport has so far centered on the charge structure factor, we note that equilibrium charge transfer across a cut provides an equivalent probe of the charge diffusion constant PhysRevLett.131.210402; wienand2023emergence: the asymptotic growth of the charge-transfer variance obeys suppmat $\langle\Delta Q(t)^{2}\rangle\simeq 2\chi\sqrt{Dt/\pi}$. With this in mind, we briefly comment on how higher cumulants of charge transfer cross over as energy and momentum relaxation terms are added. For standard diffusion, these higher cumulants scale with the usual macroscopic fluctuation form $t^{1/2}$ DerridaConjecture; Roche_2005; Bertini_2015; PhysRevLett.131.210402, while for particle–hole symmetric diffusion the $n$-th cumulant scales as $t^{n/4}$ Krajnik2022a; Krajnik2022b; Krajnik2022c; pnas.2403327121; PhysRevE.111.015410; PhysRevE.111.024141; yoshimura2025hydrodynamicfluctuationsstochasticcharged. This enhancement of charge-transfer fluctuations reflects the spatiotemporal correlations in the effective noise “bath” generated by sound waves. When energy and momentum relaxation are both present, one generically expects a crossover to the standard $t^{1/2}$ scaling, on a timescale set by the slower of the two relaxation rates. When only energy is conserved and $D_{\rm Fick}=0$, the higher moments remain anomalous at all times, as computed in Refs. pnas.2403327121; PhysRevE.111.015410; PhysRevE.111.024141; yoshimura2025hydrodynamicfluctuationsstochasticcharged. The dynamics of higher moments in the complementary limit where only momentum is conserved is more subtle, and we defer a detailed analysis to future work.
@@
Figure 2: Singular charge diffusion at weak symmetry breaking. Numerical results from the stochastic gas model in the regime of weak energy and momentum relaxation. (Left) Time-dependent charge diffusion constant $D(t)$ extracted from equilibrium current fluctuations (see suppmat), shown versus rescaled time $t/\min(\gamma_{e},\gamma_{p})$ for two fixed ratios $r\equiv\gamma_{e}/\gamma_{p}$: for $r=4.25$ (red) we show $\gamma_{e}=0.006,0.009,0.012$ (from dark to light); for $r=0.04$ (blue) we show $\gamma_{e}=0.015,0.0,0.03$ (dark to light). The black curve corresponds to the fully conserved case $\gamma_{e}=\gamma_{p}=0$ (for comparison to the weak dissipation data, we rescale time axis by 0.02t0.02\,t). The plateau value depends nontrivially on the approach direction to $(\gamma_{e},\gamma_{p})=(0,0)$, and can be either enhanced or suppressed relative to the fully conserved case. These simulations used systems of length $L=2000$. (Upper right) Total-current autocorrelator $L^{-1}\langle J(t)J(0)\rangle$ for $r=2$ (red), $r=0.02$ (blue), and fully conserved case $\gamma_{e}=\gamma_{p}=0$ (black), illustrating how the sign structure of current correlations (correlated versus anticorrelated) depends on the relative strength of the two relaxation channels. These simulations used $L=1000$. (Lower right) Weak-dissipation diffusion constant $D(r)$ obtained from the long-time plateau, plotted as a function of $r=\sqrt{\gamma_{e}/\gamma_{p}}$, showing agreement with the predicted square-root singularity in Eq. (9). These simulations used $L=2000$ and simulation times of $t=1000$.

### v2

# Noncommuting zero-noise and zero-frequency limits in particle-hole symmetric fluids

@@ Abstract
Hydrodynamics dictates how systems relax to equilibrium under generic classical or quantum dynamics Landau and Lifshitz (1987); Chaikin and Lubensky (1995); Forster (2018); Spohn (2012); Liu and Glorioso (2018). The predictions of hydrodynamics depend only on the symmetries of the system; long-lived deviations from equilibrium are due to fluctuations of the associated conserved charge densities. Within hydrodynamics, charge fluctuations generically relax diffusively (if momentum is not conserved) or spread ballistically (if it is). In the momentum-conserving case, charge density fluctuations are carried by ballistically propagating sound waves. It has long been appreciated that discrete symmetries can change this picture: for example, charge conjugation (or particle-hole) symmetry can prevent sound waves from coupling to a particular conserved charge Damle and Sachdev (1998); Fong and Schwab (2012); Lucas and Chung Fong (2018); Gopalakrishnan and Vasseur (2023). In this situation—which occurs most famously in graphene at charge neutrality Crossno et al. (2016); Majumdar et al. (2025)—ballistic energy transport coexists with diffusive charge transport. This symmetry-protected charge diffusion qualitatively differs from the “regular” hydrodynamic diffusion that occurs in the presence of momentum relaxation: as we recently showed Gopalakrishnan et al. (2024a); McCulloch et al. (2025); Gopalakrishnan et al. (2024b), symmetry-protected diffusion (in quasi-1D geometries) leads to anomalous noise signatures, such as non-gaussian full counting statistics Levitov et al. (1996).
@@
We will describe these terms in detail below. The contribution $D_{\rm Fick}$ is regular as a function of $\gamma_{e},\gamma_{p}$, with (generically) a nonzero value when $\gamma_{e}=\gamma_{p}=0$. However, at the symmetric point, there is an additional convective contribution $D_{\rm conv}$ from the ballistic sound waves to the charge diffusion constant Gopalakrishnan et al. (2018); Medenjak et al. (2020); Gopalakrishnan et al. (2024a); McCulloch et al. (2025); Hübner et al. (2025). In the momentum and energy-breaking case, the diffusion constant instead scales as
@@
with $\ldots$ representing corrections that are analytic in $\gamma_{e},\gamma_{p}$ (and disappearing as $\gamma_{e},\gamma_{p}\to 0$2. This result has important implications for recent computational approaches (on both classical and quantum computers) that aim to compute diffusion constants by zero-noise extrapolation Temme et al. (2016); Li and Benjamin (2017); Cai et al. (2023); von Keyserlingk et al. (2022); Schuster et al. (2025): the system we are considering presents an explicit example where the diffusion constant converges to some value at low noise, but this converged value is unrelated to the true zero-noise diffusion constant $\mathfrak{D}$. Previous work had established a discontinuity of this type in the fine-tuned case of integrable spin chains Nardis et al. (2022) subject to dephasing noise; we show here that the conclusion is in fact completely general, relying neither on integrability nor on the noise model.
Another key result of our work is that the diffusion constant diverges when $\gamma_{p}=0$ and $\gamma_{e}\neq 0$: i.e., when only momentum and charge are conserved, and particle-hole symmetry forbids them from coupling linearly to one another. Thus, weak momentum-conserving noise (e.g., global fluctuations of the Hamiltonian parameters) can cause a singular enhancement of charge diffusion, leading to superdiffusive charge transport. The coupled momentum and charge dynamics can be described using a recent hydrodynamic theory De Nardis et al. (2023); Roy et al. (2024b) of coupled Burgers equations. This theory was introduced to describe certain fine-tuned limits of coupled asymmetric exclusion processes; we show how it emerges without fine-tuning.
Three-mode hydrodynamics.—We briefly review the three-mode hydrodynamics of particle-hole symmetric fluids in quasi-one-dimensional geometries Gopalakrishnan et al. (2024a); McCulloch et al. (2025); Yoshimura and Krajnik (2025a, b). In this work we will be concerned with linear response about equilibrium states at nonzero temperature. We consider systems with three conserved densities: energy $e$, momentum $p$, and charge $n$. These transform under discrete symmetries as follows: $e,n$ are even under time-reversal and spatial reflection, while $p$ is odd under these symmetries; $e,p$ are even under charge conjugation, while $n$ is odd under this symmetry. (The chemical potential conjugate to $n$ is pinned by charge conjugation symmetry.) Given these constraints, we arrive at the following hydrodynamic equations Gopalakrishnan et al. (2024a):
@@
In these equations, $\eta$ denotes Gaussian white noise whose strength is set by the fluctuation-dissipation theorem, $c$ is the speed of sound, and the coefficients $C_{2},D_{\rm Fick}$ are $O(1)$ nonuniversal constants. For convenience we set $C_{2}=1$. Also, $\ldots$ denotes terms of higher order in the gradient expansion, which can be neglected for the purposes of understanding the hydrodynamics of charge Gopalakrishnan et al. (2024a). Among these neglected terms is the renormalization of the speed of sound $c$ by equilibrium charge fluctuations. This renormalization can be absorbed into the definition of $c$, so that the equations for $e,p$ form a closed system describing sound waves that that carry energy and momentum (but not charge).
@@
We will now compute this diffusion constant in two steps. First, we will artificially set $D_{\rm Fick}=0$. In this limit, the density obeys the first-order differential equation $\partial_{t}n+\partial_{x}(pn)=0$. This equation can be solved by dividing the density into small packets, each with an evolving position: $n(x,t)=\int dx_{0}\,n(x_{0},0)\,\delta(x-X(x_{0},t))$, with characteristics solving $\partial_{t}X=p(X(x_{0},t),t)$. A crucial simplification is that the ballistic sound waves move parametrically faster than the packets of charge; therefore, one can replace $p(X,t)$ with $p(x_{0},t)$ up to subleading corrections Gopalakrishnan et al. (2024a). Making this approximation, one arrives at a rigid shift of the charge pattern,
@@
To complete the calculation, we restore Fickian processes, as follows. Consider conditioning on a particular history of sound waves and a particular initial charge configuration. On this history, the charge pattern is rigidly translated by $X(t)$. Beyond this rigid translation, the charge dynamics is due to Fickian processes. If one treats the elements of charge as particles, one can describe these two effects as follows: the particle configuration is randomly rearranged over the scale $\sqrt{D_{\rm Fick}t}$ in addition to being shifted by $\sqrt{D_{\rm conv}t}$. The full diffusion constant comes from convolving these effects Gopalakrishnan et al. (2024a); Yoshimura and Krajnik (2025b), $n(x,t)=\int dx^{\prime}G_{\rm Fick}(x-x^{\prime},t)n_{0}(x^{\prime}-X(t))$: for each fixed $X$, the momentum-space charge structure factor at momentum $k$ is $S_{X}(k,t)=\chi\exp[-D_{\mathrm{Fick}}k^{2}t+ikX(t)]$; averaging this structure factor over the Gaussian variable $X(t)$ gives the final result $S(k,t)=\chi\exp[-(D_{\mathrm{Fick}}+D_{\mathrm{conv}})k^{2}t]$, which yields Eq. (1). Since the convective and Fickian effects can be treated separately, we will focus below on the fate of the convective contribution when relaxation processes are allowed; the Fickian contribution can be restored at the end of the calculation.
@@
where $S_{pp}(k,\omega)$ is the Fourier transform of $\langle p(x,t)p(0,0)\rangle$ (see 2). The diffusive broadening of the sound peaks is subleading to the Euler-scale advection in the weak symmetry breaking limit ($\gamma_{e},\gamma_{p}\to 0$). For simplicity, we now set $D_{p}=0$, (see the supplemental materials 2 for the general case). Fourier transforming Eqs. (6) and solving for $p(k,\omega)$ yields
@@
where $\chi_{e}$ and $\chi_{p}$ are energy and momentum static susceptibilities. This expression has two immediate and striking consequences. First, the diffusion constant along a general ray in $(\gamma_{e},\gamma_{p})$ space converges to a finite ray-dependent value that does not match continuously to $\mathfrak{D}$; thus the zero-noise limit is singular. Second, it suggests that when energy (momentum) remains as a residual conserved quantity, the diffusion constant vanishes (diverges). This approach can be generalized to compute the charge structure factor and optical conductivity 2. In the low frequency regime, fluctuating hydrodynamics predicts
$\sigma(\omega)=\chi\int\frac{dk}{2\pi}S_{pp}(k,\omega).$ (10)
In the d.c. limit, we recover $\sigma=\chi D$ with $D$ given by eq. (9). As illustrated in Fig. 2, the zero-noise and zero-relaxation rate results do not commute. The discontinuity of the d.c. conductivity is associated with broad regimes of anomalous super- or subdiffusive charge transport.
To gain an intuition for Eqs. (9), (10), we now consider these two limiting cases, for which the analysis simplifies.
Figure 2: Charge optical conductivity. The charge optical conductivity $\sigma(\omega)$ at weak energy and momentum relaxation, predicted by fluctuating hydrodynamics, Eq. (10). For convenience, we have set $D_{\textrm{F}ick}=0$, and $c=\chi=\chi_{e}=\chi_{p}=D_{p}=1$. (Main) When both energy and momentum conservation are broken, the conductivity has a $\gamma_{e}/\gamma_{p}$ dependent zero frequency limit. We show $\sigma(\omega)$ for several values of $\gamma_{e}$ and $\gamma_{p}$ at fixed ratio $\gamma_{p}/\gamma_{e}=10$ (red) and $\gamma_{p}/\gamma_{e}=0.1$ (blue). As the relaxation rates are reduced at fixed ratio (from light to dark), the crossover to a small frequency plateau shifts to smaller frequencies, reflecting the noncommuting zero noise and zero frequency limits. (Inset) When only momentum is relaxed, $\gamma_{e}=0$, the convective contribution to the conductivity vanishes as $\sigma(\omega)\sim\sqrt{\omega}$ in the zero frequency limit (red). Whereas, when only energy is relaxed, $\gamma_{p}=0)$, the conductivity diverges as $\sigma(\omega)\sim 1/\sqrt{\omega}$ (blue).
@@
Energy-conserving limit.—First, we consider setting $\gamma_{e}=0$, for example by considering static disorder instead of noise. In this case, energy evolves diffusively, with a diffusion constant set by $D_{e}\sim 1/\gamma_{p}$. By Lorentz covariance, momentum density is equal to the energy current: consequently, $X(t)$ is the integrated current of a diffusive quantity across a fixed point, and as such scales as $X(t)\sim t^{1/4}$. Thus, the convective contribution to charge transport becomes subdiffusive with dynamical exponent $z=4$ for any nonzero $\gamma_{p}$ when $\gamma_{e}=0$. In this regime, the convective contribution to charge transport is subleading to the Fickian term, and the diffusion constant is simply $D_{\rm Fick}$ as $\gamma_{p}\to 0$. Intuitively, the subdiffusive behavior can be understood as follows (see also Singh et al. (2021); Nardis et al. (2022)). In time $t$, a given packet of charge is repeatedly encountered by the same sound waves as they scatter back and forth. Since the backscattering of sound waves conserved energy, its momentum changes sign at each backscatter, so that successive encounters with a wave give opposite kicks. As a result, each sound wave contributes only an $O(1)$ random kick to the net displacement of a charge packet over time $t$. Since the sound waves themselves diffuse, only $O(t^{1/2})$ independent sound waves cross the origin in time $t$. As a sum of $t^{1/2}$ random kicks, the net displacement grows as $X(t)\sim t^{1/4}$. This limit clearly illustrates the crucial difference between the $\gamma_{p}=0$ and $\gamma_{p}\neq 0$ cases: in the former case each charge packet encounters each sound wave exactly once (by hydrodynamic decoupling), but for any nonzero $\gamma_{p}$ the charge packet encounters each sound wave infinitely many times (in the quasi-1D geometry we are considering).
@@
Eq. (Noncommuting zero-noise and zero-frequency limits in particle-hole symmetric fluids) describes a “degenerate” limit of two coupled degenerate Burgers equations, recently introduced in Ref. De Nardis et al. (2023) (see also Ertaş and Kardar (1992)) and explored further in Refs. Roy et al. (2024a, 2025) (where the equations take “cyclic” form yielding a Gaussian stationary state). Previously considered realizations of Eq. (Noncommuting zero-noise and zero-frequency limits in particle-hole symmetric fluids) required fine-tuning to achieve the degenerate and cyclic conditions; in the present case, these conditions follow from the symmetries of the problem. Momentum and charge are strongly nonlinearly coupled, and both quantities scale with the Kardar-Parisi-Zhang (KPZ) Kardar et al. (1986) dynamical exponent $x\sim t^{2/3}$.
@@
In the absence of relaxation processes, $\langle p(x,t)p(x,0)\rangle$ decorrelates rapidly, since sound waves move ballistically. Energy relaxation without momentum relaxation causes positive correlations, $\langle p(x,t)p(0,0)\rangle\sim(\gamma_{e}t)^{-1/2}$ on timescales starting with the first back-scattering event $\sim\gamma_{e}^{-1}$ (since momentum is now a diffusive mode with a large return probability). Conversely, momentum relaxation without energy relaxation causes sound waves to back-scatter, giving rise to anticorrelations in $\langle p(x,t)p(x,0)\rangle\sim-\gamma_{p}^{-1/2}t^{-3/2}$ (since a returning sound wave is carrying the same energy but opposite momentum) Mukerjee et al. (2006). The slower of the two relaxation processes acts as a long-time cutoff: beyond that timescale, sound waves cease to be defined 2, and $n$ is the only surviving hydrodynamic mode. For example, when energy relaxation is faster, the power-law dependence of $\langle p(x,t)p(0,0)\rangle$ is cut off on a timescale $\sim\gamma_{p}^{-1}$, so the net contribution to the diffusion constant is $\sqrt{\gamma_{p}/\gamma_{e}}$. Even though the current autocorrelation function is always instantaneously close to its value in the absence of relaxation processes, the emergent long timescale gives rise to a singular correction to diffusion (Fig. 1). We dub this phenomenon “hydrodynamic recoupling”, as the sound waves that decoupled in the $\gamma_{e}=\gamma_{n}=0$ limit “recouple” in a singular way acting as a non-Markovian bath when $\gamma_{e},\gamma_{n}>0$.
@@
On the diagonal ray $\gamma_{e}=\gamma_{p}$ backscattering of sound waves is equally likely to reverse the wave’s momentum as not, so that consecutive collisions with a charge element provide uncorrelated kicks. Therefore, charge diffusion in the weak dissipation limit along this special ray corresponds to the fully conserved case—in both cases the kicks contributing to $X(t)$ are effectively Markovian 2—fixing the coefficient of $\sqrt{\gamma_{e}/\gamma_{p}}$ to $D_{\rm conv}$ in Eq. (2). Finally we remark that the scaling form (9) will be modified for $\gamma_{p}/\gamma_{e}\ll 1$, because the nonlinearities in Eq. (Noncommuting zero-noise and zero-frequency limits in particle-hole symmetric fluids) must be included in the solution of Eq. (6).
@@
Charge transfer statistics.—While our discussion of charge transport has so far centered on the charge structure factor, we note that equilibrium charge transfer across a cut provides an equivalent probe of the charge diffusion constant McCulloch et al. (2023); Wienand et al. (2024): the asymptotic growth of the charge-transfer variance obeys 2 $\langle\Delta Q(t)^{2}\rangle\simeq 2\chi\sqrt{Dt/\pi}$. With this in mind, we briefly comment on how higher cumulants of charge transfer cross over as energy and momentum relaxation terms are added. For standard diffusion, these higher cumulants scale with the usual macroscopic fluctuation form $t^{1/2}$ Derrida et al. (2004); Roche et al. (2005); Bertini et al. (2015); McCulloch et al. (2023), while for particle–hole symmetric diffusion the $n$-th cumulant scales as $t^{n/4}$ Krajnik et al. (2022b, a, 2024); Gopalakrishnan et al. (2024a); McCulloch et al. (2025); Yoshimura and Krajnik (2025a, b). This enhancement of charge-transfer fluctuations reflects the spatiotemporal correlations in the effective noise “bath” generated by sound waves. When energy and momentum relaxation are both present, one generically expects a crossover to the standard $t^{1/2}$ scaling, on a timescale set by the slower of the two relaxation rates. When only energy is conserved and $D_{\rm Fick}=0$, the higher moments remain anomalous at all times, as computed in Refs. Gopalakrishnan et al. (2024a); McCulloch et al. (2025); Yoshimura and Krajnik (2025a, b). The dynamics of higher moments in the complementary limit where only momentum is conserved is more subtle, and we defer a detailed analysis to future work.
@@
Figure 3: Singular charge diffusion at weak symmetry breaking. Numerical results from the stochastic gas model in the regime of weak energy and momentum relaxation. (Left) Time-dependent charge diffusion constant $D(t)$ extracted from equilibrium current fluctuations (see 2), shown versus rescaled time $t/\min(\gamma_{e},\gamma_{p})$ for two fixed ratios $r\equiv\gamma_{e}/\gamma_{p}$: for $r=4.25$ (red) we show $\gamma_{e}=0.006,0.009,0.012$ (from dark to light); for $r=0.04$ (blue) we show $\gamma_{e}=0.015,0.0,0.03$ (dark to light). The black curve corresponds to the fully conserved case $\gamma_{e}=\gamma_{p}=0$ (for comparison to the weak dissipation data, we rescale time axis by 0.02t0.02\,t). The plateau value depends nontrivially on the approach direction to $(\gamma_{e},\gamma_{p})=(0,0)$, and can be either enhanced or suppressed relative to the fully conserved case. These simulations used systems of length $L=2000$. (Upper right) Total-current autocorrelator $L^{-1}\langle J(t)J(0)\rangle$ for $r=2$ (red), $r=0.02$ (blue), and fully conserved case $\gamma_{e}=\gamma_{p}=0$ (black), illustrating how the sign structure of current correlations (correlated versus anticorrelated) depends on the relative strength of the two relaxation channels. These simulations used $L=1000$. (Lower right) Weak-dissipation diffusion constant $D(r)$ obtained from the long-time plateau, plotted as a function of $r=\sqrt{\gamma_{e}/\gamma_{p}}$, showing agreement with the predicted square-root singularity in Eq. (9). These simulations used $L=2000$ and simulation times of $t=1000$.